# OmniFlatten: An End-to-end GPT Model for Seamless Voice Conversation

### v1

@@ 4.2 Training and Inference Setup
We use QWen2-0.5B Yang et al. (2024) as the base model. During the modality alignment training phase, the maximum sequence length is set to 1024 tokens. In the dialogue learning phase, the maximum sequence length is extended to 8192 tokens. We use the standard cross-entropy loss as our training objective in all stages. Additionally, during the dialogue learning phase, we apply loss masking on the User channel, as we observe that this operation enhances the stability of model training, probably due to the presence of noisy audio input in the User channel. We employ the AdamW optimizer with a weight decay of 0.1, of 0.9, and of 0.95. The maximum learning rate is set to 2e-05, with warm-up and cosine-decay. We train the model with five epochs, with the best model selected based on loss on the validation set. The batch size is set to 100 million tokens. Our code implementation is based on the NanoGPT project 444https://github.com/karpathy/nanoGPT/tree/master.
During inference, to obtain the Assistant textual response prediction from the model, we use the ground truth User channel speech in the test set as the fixed speech input and alternately fill in the predicted Assistant speech and text with fixed speech chunk and text chunk sizes.
| Librispeech(CER) | WenetSpeech(CER) | ||
| Model | test_clean | test_other | test_meeting |
| ASR | |||
| OmniFlatten (Ours) | 9.46 | 22.48 | 31.76 |
| Whisper V3 | 3.71 | 5.74 | 19.91 |
| TTS | |||
| OmniFlatten (Ours) | 10.9 | 12.87 | 50.56 |
| GT Speech Tokens | 5.82 | 12.74 | 40.18 |
@@ 4.3 Evaluations
The performance of ASR and TTS tasks after Modality Alignment
@@
The Modality Alignment training stage (Section 3.2) aims to help the model learn the correspondence between speech tokens and text tokens and acquire ASR and TTS capabilities; hence, we evaluate the effectiveness of this training stage by evaluating the ASR and TTS performance of the resulting aligned multimodal model. For the ASR evaluation, we use the model to decode the discrete speech tokens corresponding to the input speech into text output. For the TTS evaluation, we use the model to generate speech tokens based on the input text, which are then synthesized into audio using a random English female voice from CosyVoice. The synthesized audio is subsequently recognized using the Whisper Large V3 model Radford et al. (2023) 555https://huggingface.co/openai/whisper-large-v3 and the ASR output is scored against the input text. Both ASR and TTS evaluations are conducted on the publicly available Librispeech and WenetSpeech Zhang et al. (2022) datasets, and employ Character Error Rate (CER) as the evaluation metric. Note that CER could measure synthesis accuracy and robustness of a model’s TTS capabilities and also to a significant extent, reflect the audio quality. Also, the main goal of this work is learning conversation dynamics of full-duplex voice chat; therefore, for TTS evaluation in this paper, we have not adopted standard evaluation metrics for speech quality yet, such as Mean Opinion Score (MOS).
@@
For ASR evaluation, we compare the speech-text aligned multimodal model after the Modality Alignment training stage (denoted by OmniFlatten) with the Whisper Large V3 model. For TTS evaluation, GT Speech Tokens denotes discretizing the ground truth waveforms into speech tokens and detokenizing them into speech using the same English female voice. As shown in Table 1, OmniFlatten demonstrates considerable performance in both ASR and TTS tasks. These results indicate that the Modality Alignment training stage effectively transitions the unimodality text-based LLM to speech-text multimodal model with reasonable speech comprehension and generation capabilities for the following dialogue learning.
@@
The Impact of Modality Alignment and Half-duplex Dialogue Learning on Full-duplex Conversation Capability
@@
As described in Section 3.3.2, the training stage of full-duplex dialogue learning on three-stream data helps the model acquire full-duplex dialogue capacity and this model generates both speech and text for Assistant. Prior works demonstrate that competitive text-based LLMs can be reliable evaluators for various natural language generation tasks, as scores assigned by LLM evaluators on generated text show high correlations with human evaluations Zheng et al. (2023a). Therefore, we evaluate the full-duplex dialogue capabilities of OmniFlatten, by prompting a competitive text LLM to evaluate the semantics of dialogues and assign a score on the predicted Assistant textual response from the model after training on three-stream data. Notably, OmniFlatten after the final stage of training on two-stream data only outputs Assistant speech, which is challenging to be evaluated by text-based LLMs.
@@
The scoring mechanism involves designing a specific prompt and utilizing a competitive text LLM, QWen-max model 666https://help.aliyun.com/zh/model-studio/developer-reference/use-qwen-by-calling-api, to rate the responses from the model on a scale of 1 to 10 points. The specific prompt we use for LLM scoring is detailed in Appendix A. We carefully design the prompt to evaluate fluency and coherence of the predicted Assistant textual response. We also report the CE loss of the model on the test set.
@@
To analyze the impact of the modality alignment training stage (Section 3.2) and the half-duplex dialogue learning stage (Section 3.3) on the full-duplex conversation capability of OmniFlatten after training on three-stream data, we compare the LLM scores on the predicted Assistant textual responses from the following models:
QWen2-0.5B trained directly on three-stream data (denoted by OmniFlatten w/o modality alignment w/o half-duplex training).
QWen2-0.5B trained with modality alignment and full-duplex dialogue training on three-stream data (denoted by OmniFlatten w/o half-duplex training).
QWen2-0.5B trained with modality alignment, half-duplex dialogue training, and full-duplex dialogue training on three-stream data (denoted by OmniFlatten).
Ground truth textual response in the test set (denoted by GT Response).
| Model | Test Set Loss | LLM Score |
| OmniFlatten | 0.8125 | 5.185258 |
| OmniFlatten w/o half-duplex training | 0.8129 | 5.008698 |
| OmniFlatten w/o modality alignment | 0.8496 | 4.346218 |
| w/o half-duplex training | ||
| GT Response | - | 7.30685 |
| 5 | 29.2/59.4/67.4/71.9 | 3.23/129 | 2.1/5.7/8.1/17.0 | 20.55/822 | |||||||||
| 10 | 19.8/55.7/71.3/75.5 | 3.99/160 | 5.5/13.4/19.8/30.0 | 20.13/805 |
The results in Table 2 show that both Modality Alignment and Half-duplex Training stages improve the LLM scores of the predicted Assistant textual response, indicating that both stages contribute to enhancing the full-duplex dialogue performance of the model and the multi-stage training strategy effectively augments the end-to-end full-duplex spoken dialogue capabilities of the model.
Turn-taking Performance and Runtime Efficiency
To evaluate the naturalness of OmniFlatten’s full-duplex interactions, we assess whether Assistant can promptly respond after User finishes speaking (that is, Assistant taking turn), and whether Assistant can stop speaking promptly when User attemps to interrupt (that is, User taking turn). We define the following metrics.
@@
Assistant Turn-taking Acc@k: This metric is defined as whether Assistant correctly predicts a non-silence token at the k-th token after the end of a semantically meaningful speech token from User, indicating that Assistant has taken the turn and starts speaking.
@@
User Turn-taking Acc@k: This metric is defined as whether Assistant correctly outputs a silence token at the k-th token after a semantically meaningful speech token is input from User while Assistant is speaking. This metric indicates that Assistant has successfully responded to the turn-taking attempt from User by stopping Assistant’s speech and entering the state of listening. Note that in our simulated pseudo full-duplex dialogue data, User inputs are always considered as User taking turn since backchannels are not accounted for in this scenario.
The evaluation results are shown in Table 3. We make the following observations. (1) With speech chunk size 10 as used in this work, Assistant can quickly respond when User finishes speaking. Assistant Turn-taking Acc rises to 55.7% at the 5-th token and 71.3% at the 10-th token. In contrast, User Turn-taking Accs are quite poor, only 30% at the 25-th token. This is because our synthetic full-duplex data is constructed based on turn-based textual dialogues and do not cover natural scenarios of User interrupting Assistant’s speech and taking turn. In future work, we plan to refine the data synthesis pipeline and better simulate the complex interaction patterns in real-world full-duplex interactions, such as User interrupting and taking turn, and backchannels. (2) The average response time of Assistant taking turn is 160 ms and that of User taking turn is 805 ms. This discrepancy can be attributed to the fact that Assistant’s turn-taking occurs at the end of User’s turn; hence, a substantial amount of semantic information is available at that point so the system could respond promptly. Conversely, User’s turn-taking occurs at the beginning of the turn, before the semantic context is to be fully established, hence the system requires longer response time to make the decision, stopping speaking and yielding the turn. (3) We observe that using a larger speech chunk size achieves better User turn-taking accuracy and also Assistant turn-taking accuracy at large Ks. We hypothesize that this is because larger speech chunk sizes can provide more comprehensive semantic information for turn-taking prediction. However, larger chunk sizes also require longer prediction time and in turn increase Assistant’s turn-taking response time. The impact from different speech chunk sizes on User turn-taking response time is relatively minor because as explained above, User’s turn-taking is notably slow, regardless of speech chunk sizes.
5 Conclusion and Future Work
@@
In this paper, we introduce an end-to-end full-duplex spoken dialogue model OmniFlatten based on synthesizing full-duplex spoken dialogue data and designing a multi-stage progressive training paradigm for modality alignment and dialogue learning. Our approach offers a straightforward full-duplex modeling scheme as it does not change the architecture of the backbone text-based LLM nor relies on computationally intensive pre-training. Empirical evaluations show that the proposed approach is promising for developing end-to-end models to handle full-duplex interactions.
In future work, we plan to refine the data synthesis pipeline and better simulate the complex interaction patterns, such as User interrupting and taking turn, and backchannels, in real-world full-duplex interactions. Additionally, we will explore more potentials of this modeling scheme by extending to full-duplex interaction involving more modalities, such as vision.

### v2

@@ 4.2 Training and Inference Setup
We use Qwen2-0.5B Yang et al. (2024) as the base model. During the modality alignment training phase, the maximum sequence length is set to 1024 tokens. In the dialogue learning phase, the maximum sequence length is extended to 8192 tokens. We use the standard cross-entropy loss as our training objective in all stages. Additionally, during the dialogue learning phase, we apply loss masking on the user channel, as we observe that this operation enhances the stability of model training, probably due to the presence of noisy audio input in the user channel. We employ the AdamW optimizer with a weight decay of 0.1, of 0.9, and of 0.95. The maximum learning rate is set to 2e-05, with warm-up and cosine-decay. We train the model with 5 epochs, with the best model selected based on loss on the validation set. Each batch contains 100M tokens.
During inference, to obtain the assistant textual response prediction from the model, we use the ground truth user channel speech in the test set as the fixed speech input and alternately fill in the predicted assistant speech and text with the fixed speech chunk and text chunk sizes.
Model | Librispeech | Wenetspeech | ||
test-clean | test-other | test-meeting | test-net | |
Whisper-S | 3.13 | 7.37 | 25.62 | 16.66 |
Whisper-L | 1.82 | 3.5 | 18.87 | 10.48 |
VITA | 8.14 | 18.4 | 12.15 | 16.53 |
OmniFlatten | 7.91 | 19.21 | 26.1 | 19.0 |
Model | LibriTTS(WER) | AIShell-3(CER) |
Original | 2.66 | 2.52 |
ChatTTS | 8.32 | 3.87 |
CosyVoice | 2.89 | 3.82 |
OmniFlatten | 4.51 | 4.46 |
@@ 4.3 Evaluations
Evaluation of ASR and TTS Performance after Modality Alignment.
The Modality Alignment training stage (Section 3.2) aims to help the model learn the correspondence between speech tokens and text tokens and acquire ASR and TTS capabilities; hence, we evaluate the effectiveness of this training stage by evaluating the ASR and TTS performance of the multimodal-aligned model.
@@
For TTS evaluation, we use the model to generate speech tokens based on the input text, which are then synthesized into audio. We follow the experimental setup outlined in the CosyVoice paper: the synthesized audio is subsequently recognized using the Whisper-Large-V3 model Radford et al. (2023) 444https://huggingface.co/openai/whisper-large-v3 and Paraformer-zh model Gao et al. (2023) 555https://huggingface.co/funasr/paraformer-zh, for English and Chinese datasets, respectively, and the resulting ASR transcripts are scored against the input text to compute Word Error Rate (WER) for English datasets and Character Error Rate (CER) for Chinese datasets. WER/CER could measure the synthesis accuracy and robustness of a model’s TTS capabilities, and also reflect the audio quality to a significant extent. Since the main goal of this work is learning conversation dynamics of full-duplex voice chat, we have not adopted standard evaluation metrics for evaluating TTS speech quality, such as Mean Opinion Score (MOS).
@@
For ASR evaluation, we discretize the input speech into discrete speech tokens, then use the model to decode the speech tokens into text. We compare the speech-text aligned multimodal model after Modality Alignment (denoted by OmniFlatten) with Whisper-small 666https://huggingface.co/openai/whisper-small and Whisper-Large-V3 models. In addition, we have conducted a comparison of ASR capabilities with those reported by VITA. For TTS evaluation, We conduct comparisons between our approach and two competitive TTS systems, namely, ChatTTS777https://github.com/2noise/ChatTTS and CosyVoice. As shown in Table 1 and 2, OmniFlatten demonstrates considerable performance in both ASR and TTS tasks. These results indicate that the Modality Alignment training stage effectively transitions the unimodality text-based LLM to speech-text multimodal model with reasonable speech comprehension and generation capabilities for the following dialogue learning.
@@
Evaluation of Full-duplex Conversation Capability after Modality Alignment, Half-duplex Dialogue Learning, and Full-duplex Dialogue Learning.
@@
As described in Section 3.3.2, the training stage of full-duplex dialogue learning on three-stream data helps the model acquire full-duplex dialogue capacity and this model generates both speech and text. Prior works Zheng et al. (2023a) demonstrate that competitive text-based LLMs can be reliable evaluators for various natural language generation tasks, as scores assigned by LLM evaluators on generated text show high correlations with human evaluations. Therefore, we evaluate the full-duplex dialogue capabilities of OmniFlatten, by prompting a competitive text LLM to evaluate the semantics of dialogues and assign a score on the predicted assistant response from the model after training on three-stream data. We also compared three models: LLaMA-Omni, Moshi, and GLM-Voice in both Chinese and English chat capabilities. To investigate the effects of adding audio modality on the chat capability of models, we examined the text-only chat performance of Qwen2-0.5B-Instruct and Qwen2-7B-Instruct.
@@
The scoring mechanism involves designing a specific prompt and utilizing a competitive text LLM, QWen-max model 888https://help.aliyun.com/zh/model-studio/developer-reference/use-qwen-by-calling-api, to rate the responses from the model on a scale of 1 to 10 points. The specific prompt we use for LLM scoring is detailed in Appendix A. We carefully design the prompt to evaluate fluency and coherence of the predicted assistant response. Given that the full-duplex model may be disrupted by user speech in multi-turn dialogues, which could affect the assessment of chat capabilities, we exclusively focus on assessing the single-turn chat performance of the model. Given the dual output of speech and text by the model, our evaluation encompasses scoring the text output as well as converting the speech output into text using an ASR model specific to the output language, similar to the methodology described in the TTS evaluation section. The outcomes are reported separately as LLM Score (Text) and LLM Score (Speech).
@@
To analyze the impact of the modality alignment training stage (Section 3.2) and the half-duplex dialogue learning stage (Section 3.3) on the full-duplex conversation capability of OmniFlatten, we compare the LLM scores on the predicted assistant responses from the following models: (1) Trained directly on three-stream data (OmniFlatten directly 3-stream). (2) Trained with modality alignment and full-duplex three-stream dialogue data (OmniFlatten 3-stream w/o half-duplex). (3)Trained with modality alignment, half-duplex dialogue data, and full-duplex three-stream dialogue data (OmniFlatten 3-stream full process). (4) Trained with modality alignment, half-duplex dialogue data, and full-duplex three-stream dialogue data and then on two-stream data (OmniFlatten 2-stream full process).
Model | Params | En | Zh | ||
Score(Text) | Score(Speech + ASR) | Score(Text) | Score(Speech + ASR) | ||
Qwen2-0.5B-Instruct | 0.5B | 6.75 | - | 6.98 | - |
Qwen2-7B-Instruct | 7B | 8.37 | - | 8.09 | - |
LLaMA-Omni | 8B | 6.01 | 5.50 | 4.17 | 3.89 |
Moshi | 7B | 3.92 | 3.46 | - | - |
GLM-Voice | 9B | 6.97 | 6.40 | 7.02 | 6.69 |
OmniFlatten directly 3-stream | 0.5B | 2.99 | 2.59 | 4.94 | 3.95 |
OmniFlatten 3-stream w/o half-duplex | 0.5B | 3.89 | 3.54 | 5.25 | 4.76 |
OmniFlatten 3-stream full process | 0.5B | 4.88 | 3.92 | 5.6 | 5.15 |
OmniFlatten 2-stream full process | 0.5B | - | 2.19 | - | 3.06 |
Ground Truth Response | - | 7.65 | - | 6.83 | - |
Models |
Moshi | 2.9/18.8/38.5/55.1 | 553 | 0.0/6.2/14.8/45.7 | 753 | ||||||||
OmniFlatten | 20.6/53.6/ 66.3/71.7 | 193 | 10.9/30.9/41.8/51.8 | 287 |
The results, as displayed in Table 3, reveal that the multi-stage training approach enhances performance when comparing several 3-streaming models with OmniFlatten. Specifically, the OmniFlatten directly 3-stream model, lacking speech-text alignment, scores lowest; the absence of a half-duplex stage further diminishes its chat capabilities. The 2-stream experiments show a significant performance drop when models generate only speech outputs, with these models occasionally incorporating unclear semantic content in speech endings, adversely affecting scores. When comparing models that produce both speech and text, the performance in the speech modality consistently falls short of that in the text modality, indicating a slight compromise in accuracy despite the ability to synchronize speech with text outputs.
OmniFlatten exhibits superior chat performance on the English test set compared to Moshi, which frequently resorts to counter-questioning the user instead of providing direct responses. In contrast, compared with the 7B parameter LLaMA-Omni, OmniFlatten lagging scores on the English dataset while leading in the Chinese dataset. The proficiency in Chinese for LLaMA-Omni predominantly stems from the speech-text alignment process, despite the absence of Chinese data in its dialogue learning process. Comparatively, against the 9B parameter GLM-Voice model, OmniFlatten’s indicators are relatively underperforming, a situation we attribute to potential data leakage during GLM-Voice’s training process involving our test set. Additionally, comparing with the Qwen2-0.5B-Instruct results, the inclusion of the speech modality results in a noticeable decline in chat capabilities.
Turn-taking Performance and Runtime Efficiency.
To assess the naturalness of OmniFlatten’s full-duplex interactions, we evaluate the assistant’s ability to respond promptly after the user finishes speaking and to cease speaking when the user attempts to interrupt. In real-world scenarios, timely responses by machines prompt specific human interactive behaviors. For instance, if a machine fails to answer a question quickly, a human might repeat the question. Likewise, if a machine does not stop talking when interrupted, the human may try to interrupt again. Our current testing approach involves using a predefined user speech input while setting a 1.5-second threshold. We consider it a failure of user turn-taking or assistant turn-taking if, exceeded by either, the machine fails to respond (i.e., start or stop answering) within the threshold. Based on this principle, we define the following metrics:
@@
Assistant Turn-taking Acc@k: This metric is defined as whether assistant correctly predicts a non-silence token at the k-th token after the end of a semantically meaningful speech token from user, indicating that assistant starts speaking.
@@
User Turn-taking Acc@k: This metric is defined as whether assistant correctly outputs a silence token at the k-th token after a semantically meaningful speech token is input from user while assistant is speaking. This metric indicates that assistant has successfully responded to the turn-taking attempt from user by stopping assistant’s speech and listening.
The evaluation results are shown in Table 4. We compared the response times of OmniFlatten and Moshi during interactions involving user turn-taking and assistant turn-taking scenarios. It was observed that OmniFlatten demonstrates a shorter average response time compared to Moshi in both user and assistant turn-taking contexts. However, in terms of the user turn-taking metric, although OmniFlatten leads in response time and accuracy, neither model exhibits a particularly high success rate in achieving user turn-taking within 25 tokens.
5 Conclusion
@@
In this paper, we introduce an end-to-end full-duplex spoken dialogue model OmniFlatten based on synthesizing full-duplex spoken dialogue data and designing a multi-stage progressive training paradigm for modality alignment and dialogue learning. Our approach offers a straightforward full-duplex modeling scheme as it does not change the architecture of the backbone text-based LLM nor relies on computationally intensive pre-training. Empirical evaluations show that the proposed approach is promising for developing end-to-end models to handle full-duplex interactions. In future work, we plan to refine the data synthesis pipeline and better simulate the complex interaction patterns, such as user backchannels. Additionally, we will explore more potentials of this modeling scheme by extending to full-duplex interaction involving more modalities, such as vision.
6 Limitations
We acknowledge the limitations of this paper. Our work is based on training with the qwen2-0.5B model and utilizes a relatively small scale of dialogue training data. Compared to other similar models that have been scaled up, our model exhibits considerable potential for enhancement in chat capabilities. Particularly, the response speed of our model, especially in scenarios involving user turn-taking, requires further optimization. We have identified certain conflicts between the goal of early stopping by the assistant and some training objectives in speech-text alignment. Moreover, our current model is not yet capable of handling more complex human-machine interaction phenomena such as user back-channel or even assistant back-channel, indicating that the naturalness of human-machine interaction could be further enhanced.